# VMambaMorph: a Visual Mamba-based Framework with Cross-Scan Module for Deformable 3D Image Registration

### v1

@@ Abstract
Image registration, a critical process in medical imaging, involves aligning different sets of medical imaging data into a single unified coordinate system. Deep learning networks, such as the Convolutional Neural Network (CNN)-based VoxelMorph, Vision Transformer (ViT)-based TransMorph, and State Space Model (SSM)-based MambaMorph, have demonstrated effective performance in this domain. The recent Visual State Space Model (VMamba), which incorporates a cross-scan module with SSM, has exhibited promising improvements in modeling global-range dependencies with efficient computational cost in computer vision tasks. This paper hereby introduces an exploration of VMamba with image registration, named VMambaMorph. This novel hybrid VMamba-CNN network is designed specifically for 3D image registration. Utilizing a U-shaped network architecture, VMambaMorph computes the deformation field based on target and source volumes. The VMamba-based block with 2D cross-scan module is redesigned for 3D volumetric feature processing, and a fine-grained feature extraction module is proposed for high-dimensional feature learning. We validate VMambaMorph using a public benchmark brain MR-CT registration dataset, comparing its performance against current state-of-the-art methods. The results indicate that VMambaMorph achieves competitive registration quality. The code for VMambaMorph is available on GitHub.
@@ 1 Introduction
    A fine-grained, simple, yet efficient feature extractor is utilized for high-dimensional feature learning prior to registration.
@@
Figure 1: The Framework of VMambaMorph, and the Details of Visual State Space Block.
@@ 2 Method
The proposed VMamaMorph is sketched in Figure 1, consisting of a two-branch feature extractor module, registration module, and STN. A pair of target volume and source volume are respectively denoted as $\textbf{{I}}^{\rm tgt}$ and $\textbf{{I}}^{\rm src}$, and their corresponding segmentation is denoted as $\textbf{{S}}^{\rm tgt}$ and $\textbf{{S}}^{\rm src}$. The registration module is $\mathcal{R}(\cdot,\cdot)$ and the feature extractor is $\mathcal{F}(\cdot)$. The deformation field is denoted as $\phi$. The losses include $\mathcal{L}_{\rm dice}(\cdot,\cdot)$ and $\mathcal{L}_{\rm smooth}(\cdot)$ following [1, 11].
@@ 2.2 Training Setting
$\textbf{{F}}^{\{tgt,src\}}=\mathcal{F}(\textbf{{I}}^{\{tgt,src\}};\textbf{{w}})$ (11)
@@ 3 Experiments
Dataset: A public benchmark brain MR-CT registration dataset named SR-Reg is utilized in this paper [11]. The SR-Reg dataset consists of 180 well-aligned, skull-stripped and intensity-rectified MR-CT pairs, with corresponding segmentation labels, and it is mainly based on the SynthRAD 2023 dataset [23, 2, 13]. The original size of each volume is with $192\times 208\times 176$ voxels with a resolution of $1\times 1\times 1mm^{3}$. More details can be found at  [^1] https://github.com/Guo-Stone/MambaMorph. Considering the input size requirement of VMamba, and computational cost on a single GPU, we have resized the volume to $128\times 128\times 128$, and randomly select 150 cases for training, 10 cases for validation, and the rest 20 cases for testing.
@@
VMambaMorph 78.46$\pm$2.13 1.69$\pm$0.22 0.33$\pm$0.01 2.94 9.35 0.07
VMambaMorph 82.49$\pm$1.99 1.38$\pm$0.18 0.35$\pm$0.01 3.25 9.64 0.10
[…]
@@
Ablation Study: To validate the effectiveness of the fine-grained feature extractor, we conducted an ablation study. The performance impact of the feature extractor (denoted as FeatEx) is evaluated in conjunction with all methods in Table 1. The deployment of FeatEx is indicated with a ✓, while its absence is marked with an ✗. All methods demonstrate improvement when the feature extractor is utilized.
@@ 4 Conclusion
In this paper, we introduce VMambaMorph, a multi-modality deformable image registration framework based on the Visual State Space Model. The architecture of VMambaMorph is developed with a hybrid registration module with 3D VMamba-based network blocks and convolutional network blocks. Following the recent advances in feature learning for multi-modality registration, VMambaMorph incorporates a fine-grained feature extractor to achieve high-level spatial features more effectively. Our validation on a public MR-CT registration dataset indicates that VMambaMorph surpasses recent methodologies such as MambaMorph, as well as established techniques like VoxelMorph and TransMorph. This paves the way for future applications of VMamba in broader computer vision tasks.

### v2

# VMambaMorph: a Multi-Modality Deformable Image Registration Framework based on Visual State Space Model with Cross-Scan Module

@@ Abstract
Image registration, a critical process in medical imaging, involves aligning different sets of medical imaging data into a single unified coordinate system. Deep learning networks, such as the Convolutional Neural Network (CNN)-based VoxelMorph, Vision Transformer (ViT)-based TransMorph, and State Space Model (SSM)-based MambaMorph, have demonstrated effective performance in this domain. The recent Visual State Space Model (VMamba), which incorporates a cross-scan module with SSM, has exhibited promising improvements in modeling global-range dependencies with efficient computational cost in computer vision tasks. This paper hereby introduces an exploration of VMamba with image registration, named VMambaMorph. This novel hybrid VMamba-CNN network is designed specifically for 3D image registration. Utilizing a U-shaped network architecture, VMambaMorph computes the deformation field based on target and source volumes. The VMamba-based block with 2D cross-scan module is redesigned for 3D volumetric feature processing. To overcome the complex motion and structure on multi-modality images, we further propose a fine-tune recursive registration framework. We validate VMambaMorph using a public benchmark brain MR-CT registration dataset, comparing its performance against current state-of-the-art methods. The results indicate that VMambaMorph achieves competitive registration quality. The code for VMambaMorph with all baseline methods is available on GitHub.
@@ 1 Introduction
    To overcome the challenge of complex motion and structure in multi-modality image registration, a recursive registration framework is adopted. In addition, a fine-grained feature extractor is utilized before registration.
@@
Figure 2: The Architecture of VMambaMorph, and the Details of Visual State Space Block.
@@ 2 Method
The proposed VMamaMorph-based recursive registration framework is sketched in Figure 1 and the network structure of VMambaMorph is shown in 2, consisting of a two-branch feature extractor module, registration module, and STN. A pair of target volume and source volume are respectively denoted as $\textbf{{I}}^{\rm tgt}$ and $\textbf{{I}}^{\rm src}$, and their corresponding segmentation is denoted as $\textbf{{S}}^{\rm tgt}$ and $\textbf{{S}}^{\rm src}$. The registration module is $\mathcal{R}(\cdot,\cdot)$ and the feature extractor is as $\mathcal{F}(\cdot)$. The deformation field is denoted as $\phi$. The losses include $\mathcal{L}_{\rm dice}(\cdot,\cdot)$ and $\mathcal{L}_{\rm smooth}(\cdot)$ following [1, 11].
@@ 2.1 Visual State Space Model
2.2 Recursive Registration Framework
Addressing the intricate challenge presented by complex motion and structural variations in multi-modality image registration, we adopt a recursive registration framework structure aimed at achieving coarse-to-fine registration, as depicted in Fig. 1. This initiative is inspired by [33, 37]. In our proposed method, the residual transformation $\varphi_{k}$, representing the alignment discrepancy between the target image $\textbf{{I}}^{\rm tgt}$ and the successively warped source feature map $\phi_{k-1}(\textbf{{I}}^{\rm src})$ from the previous $k-1$ registration level, is computed by the operator $\mathcal{R}$ and progressively refined through compositional accumulation:
$\left\{\begin{array}[]{cc}\phi_{k}=\phi_{k-1}\circ\varphi_{k}\\ \varphi_{k}=\mathcal{R}(\phi_{k-1}(\textbf{{I}}^{\rm src}),\textbf{{I}}^{\rm tgt})\end{array}\right.,~{}~{}0<k\leq K$ (9)
where $\circ$ signifies the composition of two deformation fields, with $\phi_{0}$ being initialized without deformation field. The proposed VMambaMorph network, as illustrated in Fig. 2, integrates a weight-sharing, enhanced fine-grained feature extractor and a U-shaped VMamba-CNN hybrid registration module to accurately estimate the deformation field $\psi_{k}$, and then to refine the final prediction $\phi_{k}$. Following the optimal setting validated by [36], the recursive number is set to ${K}_{\rm t}=2$ in the training process, and varying recursive number ${K}_{\rm i}>1$ is compared in the inference stage to optimize the registration performance.
2.3 Enhanced Fine-grained Feature Extractor
@@ 2.4 Training Setting
$\textbf{{F}}^{\{\rm tgt,src\}}=\mathcal{F}(\textbf{{I}}^{\{\rm tgt,src\}};\textbf{{w}})$ (12)
@@ 3 Experiments
Dataset: A public benchmark brain MR-CT registration dataset named SR-Reg is utilized in this paper [11]. The SR-Reg dataset consists of 180 well-aligned, skull-stripped and intensity-rectified MR-CT pairs, with corresponding segmentation labels, and it is mainly based on the SynthRAD 2023 dataset [23, 2, 13]. The original size of each volume is with $192\times 208\times 176$ voxels with a resolution of $1\times 1\times 1mm^{3}$. More details can be found at [11]. Considering the input size requirement of VMamba, and computational cost on a single GPU, we have resized the volume to $128\times 128\times 128$, and randomly select 150 cases for training, 10 cases for validation, and the rest 20 cases for testing.
@@
VMambaMorph 82.94$\pm$2.01 1.35$\pm$0.18 1.04$\pm$0.05 3.25 9.64 0.10
Metrics: We use the the percentage of mean dice-coefficients (Dice), 95% Hausdorff distance (HD95) to measure the registration accuracy, percentage of non-positive Jacobian determinant (P|J∅|≤0) to evaluate the diffeomorphic property of deformation field. The inference time (Time) with the unit of second (s), GPU memory cost (Mem) with the unit of Gigabyte (Gb) and the amount of parameters of network with the unit Mb is used to evaluate the proposed VMamabaMorph with other baseline methods.
@@
Table 2: Ablation study of VMambaMorph methods for feature extractor (FeatEx) and recursive training ($K_{\rm t}$) or inference ($K_{\rm i}$).
Model $K_{\rm t}$ $K_{\rm i}$ FeatEx Dice HD95 P|J∅|≤0 Mem Par Time
(%) vox (%) (Gb) (Mb) (sec)
Initial - - - 62.42$\pm$3.29 3.73$\pm$0.41 - - - -
VMambaMorph 1 1 78.46$\pm$2.13 1.69$\pm$0.22 0.33$\pm$0.01 2.94 9.35 0.07
VMambaMorph 1 1 82.49$\pm$1.99 1.38$\pm$0.18 0.35$\pm$0.01 3.25 9.64 0.10
VMambaMorph 1 2 82.84$\pm$2.05 1.44$\pm$0.20 1.05$\pm$0.01 3.93 9.64 0.20
VMambaMorph 1 3 82.80$\pm$2.06 1.47$\pm$0.18 1.71$\pm$0.07 3.93 9.64 0.29
VMambaMorph 1 4 82.62$\pm$2.11 1.50$\pm$0.18 2.29$\pm$0.09 3.93 9.64 0.36
VMambaMorph 2 1 82.47$\pm$2.01 1.38$\pm$0.18 0.35$\pm$0.01 3.93 9.64 0.10
VMambaMorph 2 2 82.94$\pm$2.01 1.35$\pm$0.18 1.04$\pm$0.05 3.93 9.64 0.19
VMambaMorph 2 3 82.85$\pm$2.07 1.47$\pm$0.18 1.69$\pm$0.06 3.93 9.64 0.28
VMambaMorph 2 4 82.65$\pm$2.16 1.51$\pm$0.17 2.26$\pm$0.08 3.93 9.64 0.40
@@
Ablation Study: To validate the effectiveness of the fine-grained feature extractor, we conducted an ablation study. The performance impact of the feature extractor (denoted as FeatEx) is evaluated in conjunction with all methods in Table 2. The deployment of FeatEx is indicated with a ✓, while its absence is marked with an ✗. All methods demonstrate improvement when the feature extractor is utilized. To validate the proposed recursive registration framework, the bottom of Table 2 further explores the affect of varying recursive numbers in training and inference.
@@ 4 Conclusion
In this paper, we introduce VMambaMorph, a multi-modality deformable image registration framework based on the Visual State Space Model. The architecture of VMambaMorph is developed with a hybrid registration module with 3D VMamba-based network blocks and 3D CNN-based network blocks. Following the recent advances in feature learning for multi-modality registration, VMambaMorph incorporates the recursive registration framework with a pair of fine-grained feature extractor to achieve high-level spatial features more effectively. Our validation on a public MR-CT registration dataset indicates that VMambaMorph surpasses recent methodologies such as MambaMorph, as well as established techniques like VoxelMorph and TransMorph. This paves the way for future applications of VMamba in broader computer vision tasks.